OpenStudy (anonymous):

Simplify the radical expression by rationalizing the denominator. 5/√30

OpenStudy (anonymous):

\[\frac{ 5 }{ \sqrt{30} }\]

OpenStudy (anonymous):

multiply top and bottom by \(\sqrt{30}\)

OpenStudy (anonymous):

then cancel the 5

OpenStudy (luigi0210):

\[\frac{ 5 }{ \sqrt{30} }*\frac{ \sqrt{30} }{ 30 }=\frac{ 5\sqrt{30} }{ 30 }=\frac{ \sqrt{30} }{ 6 }\]

OpenStudy (anonymous):

Thank you!

OpenStudy (luigi0210):

Yup :P

OpenStudy (anonymous):

Could you help me with one more?

OpenStudy (anonymous):

Solve the equation. Identify any extraneous solutions. √a = -2

OpenStudy (luigi0210):

we t solve for this you would square both sides..\[(\sqrt{a})^2=(-2)^2\] a=4

OpenStudy (luigi0210):

\[\sqrt{4}=\pm2\]

OpenStudy (anonymous):

Thanks!

OpenStudy (luigi0210):

You're welcome :)